OpenStudy (anonymous):

Solve U = 0.5CV2 for C. I need help!!

OpenStudy (nikato):

just dividfe both sides by 0.5V2

OpenStudy (anonymous):

here are the options

OpenStudy (anonymous):

Please, no direct answers, but instead guide. Just saying. :)

OpenStudy (yttrium):

Just divide the whole equation by 0.5V^2 and knowing that 0.5 = 1/2 So what you'll get? @acacia14

OpenStudy (anonymous):

I don't know

OpenStudy (yttrium):

To solve for C, we have to isolate all existing variable except C. So, 0.5CV^2 = U Therefore, (0.5CV^2)/0.5V^2 = U/0.5V^2 Now, tell me what would be the value of C.

OpenStudy (anonymous):

the value of c is 2

OpenStudy (yttrium):

Nope.

OpenStudy (yttrium):

C = U/0.5V^2 which is also equal to what?

OpenStudy (yttrium):

@acacia14

OpenStudy (anonymous):

C=2u/v^2

OpenStudy (yttrium):

you got this